对于事件 AA 和 BB,已知 P(A)=0.38\mathrm{P}(A) = 0.38,P(A∣B)=0.5\mathrm{P}(A \mid B) = 0.5 且 P(A∪B)=0.52\mathrm{P}(A \cup B) = 0.52。

  1. (i)

    求 P(B)\mathrm{P}(B)。

  2. (ii)

    对于第三个事件 CC,已知 P(C)=0.6\mathrm{P}(C) = 0.6 且 P((A′∪B′)∣C)=0.85\mathrm{P}((A' \cup B') \mid C) = 0.85。
    求 P(A∩B∩C)\mathrm{P}(A \cap B \cap C)。

  3. (iii)

    说明事件 AA 和 CC 是否互斥,并给出理由。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

10(i)

1. Express the probability of the intersection, $\mathrm{P}(A \cap B)$, in terms of $\mathrm{P}(B)$.

Using the conditional probability formula P(A∩B)=P(A∣B)P(B)\mathrm{P}(A \cap B) = \mathrm{P}(A \mid B)\mathrm{P}(B):
P(A∩B)=0.5P(B)\mathrm{P}(A \cap B) = 0.5\mathrm{P}(B)

2. Set up an equation for $\mathrm{P}(B)$ using the addition rule.

Using the addition rule P(A∪B)=P(A)+P(B)−P(A∩B)\mathrm{P}(A \cup B) = \mathrm{P}(A) + \mathrm{P}(B) - \mathrm{P}(A \cap B):
0.52=0.38+P(B)−0.5P(B)0.52 = 0.38 + \mathrm{P}(B) - 0.5\mathrm{P}(B)
0.52=0.38+0.5P(B)0.52 = 0.38 + 0.5\mathrm{P}(B)

3. Solve the linear equation for $\mathrm{P}(B)$.

0.5P(B)=0.52−0.380.5\mathrm{P}(B) = 0.52 - 0.38
0.5P(B)=0.140.5\mathrm{P}(B) = 0.14
P(B)=0.140.5=0.28\mathrm{P}(B) = \frac{0.14}{0.5} = 0.28

答案

0.28

10(ii)

1. Apply De Morgan's laws to simplify the event given in the conditional probability.

A′∪B′=(A∩B)′A' \cup B' = (A \cap B)'
P((A∩B)′∣C)=0.85\mathrm{P}((A \cap B)' \mid C) = 0.85

2. Use the complement rule for conditional probabilities to find $\mathrm{P}(A \cap B \mid C)$.

P(A∩B∣C)=1−P((A∩B)′∣C)\mathrm{P}(A \cap B \mid C) = 1 - \mathrm{P}((A \cap B)' \mid C)
P(A∩B∣C)=1−0.85=0.15\mathrm{P}(A \cap B \mid C) = 1 - 0.85 = 0.15

3. Apply the conditional probability formula to find the intersection of all three events.

P(A∩B∩C)=P(A∩B∣C)×P(C)\mathrm{P}(A \cap B \cap C) = \mathrm{P}(A \cap B \mid C) \times \mathrm{P}(C)
P(A∩B∩C)=0.15×0.6=0.09\mathrm{P}(A \cap B \cap C) = 0.15 \times 0.6 = 0.09

答案

0.090.09

10(iii)

1. Establish the condition for mutually exclusive events and its implication on the three-event intersection.

If events AA and CC are mutually exclusive, then P(A∩C)=0\mathrm{P}(A \cap C) = 0.
Since (A∩B∩C)(A \cap B \cap C) is a subset of (A∩C)(A \cap C), it must follow that P(A∩B∩C)=0\mathrm{P}(A \cap B \cap C) = 0 as well.

2. Compare with the result from part (ii) to make a conclusion.

From part (ii), we found that P(A∩B∩C)=0.09\mathrm{P}(A \cap B \cap C) = 0.09.
Since 0.09≠00.09 \neq 0, this implies P(A∩C)≥0.09\mathrm{P}(A \cap C) \ge 0.09, and thus P(A∩C)≠0\mathrm{P}(A \cap C) \neq 0.
Therefore, events AA and CC are not mutually exclusive.

答案

No, because P(A∩C)≥0.09\mathrm{P}(A \cap C) \ge 0.09, so P(A∩C)≠0\mathrm{P}(A \cap C) \neq 0.

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1已知事件 AA 和 BB 满足 P(B)=13\mathrm{P}(B) = \frac{1}{3},P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} 且 P(A∪B)=35\mathrm{P}(A \cup B) = \frac{3}{5}。…
  2. 2凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  3. 3一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  4. 4已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…
  5. 5对于事件 AA 和 BB,已知 P(A)=0.55\mathrm{P}(A) = 0.55,P(B)=0.65\mathrm{P}(B) = 0.65 且 P(A∪B)=0.93\mathrm{P}(A \cup B) = 0.93。 求 P(A∩B)\mathrm{P}(A \cap B)…